AIR-BROADENING AND SHIFT PARAMETERS IN THE ν_{3} BAND OF OZONE

\underline{M. A. H. SMITH}, Science Directorate, NASA Langley Research Center, Hampton, VA 23681-2199; V. MALATHY DEVI and D. CHRIS BENNER, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795.

Line parameter errors can contribute significantly to the total errors in retrievals of terrestrial atmospheric ozone concentration profiles using the strong 9.6-μm band, particularly for nadir-viewing experiments^{a}. Detailed knowledge of the interfering ozone signal is also needed for retrievals of other atmospheric species^{b} in this spectral region. We have determined Lorentz air-broadening and pressure-induced shift coefficients along with their temperature dependences for a number of transitions in the ν_{3} fundamental band of ^{16}O_{3}. These results were obtained by applying the multispectrum nonlinear least-squares fitting technique^{c} to a set of 31 high-resolution infrared absorption spectra of O_{3} previously recorded at temperatures between 160 and 300 K with several different room-temperature and coolable sample cells at the McMath-Pierce Fourier transform spectrometer at the National Solar Observatory on Kitt Peak. We compare our results with other available measurements and with the ozone line parameters in the 2008 HITRAN database^{d} .

aJ. Worden et al., J. Geophys. Res. \underline{109} (2004) 9308-9319.

bR. Beer et al., Geophys. Res. Lett. \underline{35} (2008) L09801.

cD. Chris Benner et al., JQSRT \underline{53} (1995) 705-721.

dL. S. Rothman et al., JQSRT \underline{110} (2009) 533-572.